# Mittag-Leffler Quantum Statistics and Thermodynamic Anomalies

### v1

@@ I Introduction
Several generalized statistical frameworks have been developed along these lines. Tsallis statistics replace the exponential with a power-law function derived from a nonextensive entropy, introducing a parameter $q$ that captures deviations from standard thermodynamics [1, 2, 3, 4, 5]. Kaniadakis statistics introduce a deformation parameter $\kappa$ to construct a relativistically motivated generalization of the exponential function [6]. Superstatistics takes a different approach by averaging fluctuations in intensive parameters (such as temperature), thereby modeling systems as superpositions of local equilibria [7]. Each of these frameworks distinctly extends the traditional Boltzmann factor, allowing statistical mechanics to address more complex or non-equilibrium systems.
@@
Motivated by efforts to generalize the Boltzmann factor, we employ the Mittag-Leffler (ML) function, a natural extension of the exponential function widely used to model complex systems exhibiting anomalous dynamics  [8, 9, 10]. The ML function features a tunable parameter, $\alpha$, which enables it to capture deviations from standard behavior and extend traditional statistical distributions. Building on this, in our earlier study  [11], we proposed the MLMB distribution, derived by generalizing the classical MB distribution through the expansion of the standard exponential function into the ML function. This generalization, parameterized by $\alpha$, enables a more comprehensive representation of particle interactions, encompassing behaviors not captured by the standard MB distribution. Notably, this formulation provides a continuous transition between classical and quantum statistical regimes, reproducing fermion-like and boson-like characteristics, thus effectively modeling interaction-induced deviations within a unified theoretical framework.
@@
Building upon the ML generalization of the Maxwell-Boltzmann distribution [11], the present study extends this framework to the quantum domain by introducing the Mittag-Leffler Bose-Einstein (MLBE) and Mittag-Leffler Fermi-Dirac (MLFD) distributions -ML-based distributions-. These novel distributions are formulated by generalizing the exponential function in the canonical BE and FD distributions to the ML function, thereby introducing the tunable parameter $\alpha$ into the quantum statistical weights. This extension preserves the essential quantum characteristics of bosons and fermions while allowing for a more flexible description of particle interactions and thermodynamic behavior, particularly in systems that deviate from idealized equilibrium conditions.
@@
To investigate the physical implications of these generalized distributions, we employ the framework of thermodynamic geometry, which provides a powerful geometric interpretation of thermodynamic state space. By analyzing the thermodynamic curvature and associated metrics, this approach reveals information about microscopic interactions, stability, and phase transitions within the system. Thermodynamic geometry has previously provided important information on systems governed by generalized statistics such as Tsallis, Kaniadakis, and superstatistics [12, 13, 14, 15, 16, 17, 18], and its application to the MLBE and MLFD distributions provides detailed information on their thermodynamic behavior, including heat capacity and condensation phenomena.
@@ II Broader Context of Generalized Statistics and the Role of Mittag-Leffler Distributions
where $t$ denotes time, $\tau$ is a characteristic relaxation time, and $\beta$ ($0<\beta<1$) is the stretching exponent that quantifies deviations from Debye’s exponential law. At long times, such stretched exponentials asymptotically cross over into power-law decay [19]. This crossover is precisely captured by the one-parameter ML function $E_{\alpha}(-t^{\alpha})$, where $\alpha$ ($0<\alpha\leq 1$) is the fractional order parameter: for $\alpha=1$ the function reduces to a pure exponential, while for $0<\alpha<1$ it exhibits algebraic tails.
@@
where $x$ is the displacement of a diffusing particle and the exponent $\alpha$ quantifies the deviation from normal Brownian diffusion (with $\alpha=1$ corresponding to ordinary diffusion, $\alpha<1$ to subdiffusion, and $\alpha>1$ to superdiffusion). Such scaling laws emerge directly from fractional Fokker–Planck equations whose unique solutions are expressed in terms of ML functions [20, 21].
@@
where $\mathcal{L}$ denotes the Laplace transform, $s$ is the Laplace variable conjugate to time $t$, and $\lambda>0$ is a characteristic rate parameter. This functional form arises naturally from fractional operators and cannot be represented by ordinary exponential kernels [22].
These examples illustrate that the ML function is far more than a mathematical curiosity: it is central to the physics of systems with long memory, anomalous transport, and hierarchical energy landscapes. Indeed, ML functions have long been recognized as the “queen function of fractional calculus,” emerging naturally as the fundamental solution of fractional-order differential equations that describe anomalous diffusion, non-exponential relaxation, and long-memory processes [8, 23]. Its relevance spans numerous areas of physics: modeling viscoelastic relaxation via the fractional Zener model, where stress $\sigma(t)$ and strain $\varepsilon(t)$ are related by fractional derivatives [24, 25]; describing anomalous transport and dispersion in porous or plasma environments via fractional advection–dispersion equations [26, 20]; and characterizing non-Debye dielectric relaxation using the Prabhakar (three-parameter ML) function [27]. Additionally, ML functions provide analytic solutions to fractional kinetic and reaction–diffusion equations relevant in astrophysics, space sciences, and time-series analysis [28, 29].
@@ III Mittag-Leffler Function As a Generalized Distribution
where $\Gamma$ is Gamma function and $\alpha>0$. In the limit $\alpha=1$, the function reduces to the standard exponential $E_{1}(X)=e^{X}$. Moreover, when $\alpha\gg 1$, $E_{\alpha}(X)$ asymptotically approaches 1, reflecting a uniform limiting behavior. The one-parameter form $E_{\alpha}(X)$ seamlessly interpolates between exponential and power-law behaviors. Further generalizations to multi-parameter versions, such as $E_{\alpha,\beta}(X)$ and $E_{\alpha,\beta}^{\gamma}(X)$, enhance this flexibility, enabling more accurate modeling of memory effects and anomalous diffusion in complex systems [30, 23].
Having outlined its mathematical foundation, we now recall the application of the ML function in classical statistical mechanics, where it has been used to generalize the MB distribution  [11]. We introduced the MLMB distribution by generalizing the exponential function in the classical MB distribution to the ML function. This modification enabled a smooth interpolation between classic and quantum (fermion-like and boson-like) behaviors, effectively capturing interaction-induced deviations from classical statistics. In particular, the MLMB distribution exhibits a transition between the classical behavior and the fermion-like (or boson-like) behavior depending on the value of the parameter $\alpha$. When $0<\alpha<1$, the system displays fermionic traits, including negative thermodynamic curvature indicative of repulsive interactions similar to the Pauli exclusion principle. For $1<\alpha\leq 1.5$, it shows bosonic characteristics, with attractive interactions signaling positive curvature and the emergence of Bose-Einstein condensation as the fugacity approaches a critical value. This work demonstrated the versatility of the ML function in extending classical statistical mechanics beyond standard assumptions.
@@
In the MLMB classical distribution [11], the deformation parameter $\alpha$ governs these effective interactions:
@@
This approach parallels other generalized statistical frameworks. For instance, the Tsallis entropic parameter $q$ and the Kaniadakis deformation parameter $\kappa$ [31, 32]similarly tune effective interaction strength and the sign of thermodynamic curvature. The MLMB distribution extends these ideas by enabling a continuous transition between fermion-like (repulsive) and boson-like (attractive) regimes through the single parameter $\alpha$.
@@
As shown in Fig.  LABEL:n_MLFD, the MLFD distribution yields an occupation number that remains positive for all values of $\alpha$, approaching unity as $X\to$-$\infty$ and vanishing as $X\to+\infty$. It is noteworthy that for the MLFD distribution with $\alpha>1$, the distribution exhibits anomalous behavior: in the negative domain of $X$, it temporarily exceeds unity before asymptotically approaching 1.
In contrast, Fig. LABEL:n_MLBE shows that while the MLBE distribution also satisfies $n(X)\to 0$ as $X\to+\infty$, it exhibits divergence near $X=0$ and produces negative values for $X<0$. These negative values of distribution are unphysical, as the occupation number must remain non-negative to ensure physical consistency. The MLFD distribution satisfies this criterion for all $X$ and the entire range of $\alpha$. However, the MLBE distribution imposes a constraint on the fugacity, requiring it to remain within the range $0\leq z\leq 1$ to exclude unphysical results. At the critical value $1<z$, the function $n(X)$ diverges, reflecting the characteristic behavior of the standard BE distribution, where BE condensation occurs [33, 34, 35].
@@ V Thermodynamic Curvature of the MLBE and MLFD statistics
Thermodynamic geometry, as developed notably by Ruppeiner and Weinhold, provides a differential geometric structure to the space of thermodynamic parameters [36, 37]. Within this framework, the space of equilibrium states is treated as a Riemannian manifold, where thermodynamic fluctuations and interactions can be examined through the properties of a metric tensor. The Ruppeiner metric is derived by taking the negative second derivative of entropy with respect to extensive thermodynamic variables, such as internal energy, volume, and particle number. Alternatively, in the energy representation proposed by Weinhold, the metric is obtained from the Hessian of the internal energy with respect to its natural variables.
Applying Legendre transformations to fundamental thermodynamic potentials enables us to construct the equivalent metric structures based on various potentials, such as the Helmholtz and Gibbs free energies. Furthermore, the Fisher information metric—another significant geometric tool—can be introduced through the second derivatives of the logarithm of the partition function with respect to the intensive parameters [38, 39]:
@@
In systems characterized by two thermodynamic degrees of freedom, the associated geometry is two-dimensional. Janyszek and Mrugała demonstrated that when a metric is derived directly from a thermodynamic potential, the scalar curvature of the parameters space can be expressed in terms of the potential’s second and third derivatives [40]. In two dimensions, the Ricci scalar curvature is computed via:
@@
For an ideal classical gas described by the MB statistics, the thermodynamic curvature vanishes. In contrast, quantum ideal gases exhibit non-zero curvature: it is positive for BE statistics and negative for FD statistics. These signs reflect effective statistical interactions—attractive in the case of bosons and repulsive for fermions [38, 39, 18]. The divergence of curvature near a critical fugacity (e.g., $z\to 1$ for BE) can signal the presence of phase transitions such as Bose-Einstein condensation. The theoretical framework of geometric thermodynamics has been addressed in earlier studies [41, 42, 35, 32, 43], which provide comprehensive reviews of its geometric structure and thermodynamic implications.
@@
In the MLBE distribution, for $\alpha<1$, the system displays a fugacity-dependent crossover between fermionic and bosonic characteristics. At a specific point $z=z^{*}_{\alpha}$, indicated by a solid dot in Fig. LABEL:R_MLBE_kam, the thermodynamic curvature changes sign. For fugacity values less than $z^{*}_{\alpha}$, the system displays fermionic characteristics, marked by effective repulsive interactions. In contrast, for $z>z^{*}_{\alpha}$, the system transitions into the bosonic regime, characterized by attractive effective interactions between particles. This crossover, along with the dependence of $z^{*}_{\alpha}$ on the parameter $\alpha$, is illustrated in Fig. LABEL:R_MLBE_kam.
In contrast, when $\alpha>1$, the MLBE distribution exhibits predominantly bosonic behavior for all values of $z$, as indicated by the consistently positive thermodynamic curvature shown in Fig. LABEL:R_MLBE_bish. Moreover, irrespective of the value of $\alpha$, the MLBE distribution features a critical point at $z^{\alpha}_{c}=z^{\alpha}_{c}=1$, where the thermodynamic curvature diverges. This divergence signals the onset of a phase transition analogous to Bose-Einstein condensation, as discussed in Refs. [39, 35]. The critical fugacity $z_{c}$ thus marks the emergence of macroscopic quantum phenomena associated with condensation.
@@
Turning to the MLFD distribution, a similar classification emerges based on $\alpha$. For $\alpha<1$, the system exhibits purely fermionic behavior: the thermodynamic curvature remains negative across all values of $z$, starting from strongly negative values at small $z$ and becoming less negative as $z$ increases, see Fig. (LABEL:R_MLFD_kam). In contrast, for $\alpha>1$, the thermodynamic curvature $R$ initially assumes positive values at very small values of $z$. At a specific threshold $z=z^{*}_{\alpha}$, as summarized in the accompanying table in Fig. LABEL:R_MLFD_bish, the thermodynamic curvature undergoes a sign change. Beyond this point, $R$ rapidly transitions to negative values and gradually stabilizes with a mildly decreasing slope as $z$ increases, as shown in Fig. LABEL:R_MLFD_bish.

### v2

@@ I Introduction
Several generalized statistical frameworks have been developed along these lines. Tsallis statistics replace the exponential with a power-law function derived from a nonextensive entropy, introducing a parameter $q$ that captures deviations from standard thermodynamics [jiulin2007nonextensivity, cohen2002statistics, plastino2004Tsallis, tamarit2005relaxation, Tsallis2002mixing]. Kaniadakis statistics introduce a deformation parameter $\kappa$ to construct a relativistically motivated generalization of the exponential function [kaniadakis2021new]. Superstatistics takes a different approach by averaging fluctuations in intensive parameters (such as temperature), thereby modeling systems as superpositions of local equilibria [beck2003superstatistics]. Each of these frameworks distinctly extends the traditional Boltzmann factor, allowing statistical mechanics to address more complex or non-equilibrium systems.
@@
Motivated by efforts to generalize the Boltzmann factor, we employ the Mittag-Leffler (ML) function, a natural extension of the exponential function widely used to model complex systems exhibiting anomalous dynamics  [gorenflo2020mittag, dos2020mittag, mathai2010mittag]. The ML function features a tunable parameter, $\alpha$, which enables it to capture deviations from standard behavior and extend traditional statistical distributions. Building on this, in our earlier study  [seifi2025intrinsic], we proposed the MLMB distribution, derived by generalizing the classical MB distribution through the expansion of the standard exponential function into the ML function. This generalization, parameterized by $\alpha$, enables a more comprehensive representation of particle interactions, encompassing behaviors not captured by the standard MB distribution. Notably, this formulation provides a continuous transition between classical and quantum statistical regimes, reproducing fermion-like and boson-like characteristics, thus effectively modeling interaction-induced deviations within a unified theoretical framework.
@@
Building upon the ML generalization of the Maxwell-Boltzmann distribution [seifi2025intrinsic], the present study extends this framework to the quantum domain by introducing the Mittag-Leffler Bose-Einstein (MLBE) and Mittag-Leffler Fermi-Dirac (MLFD) distributions -ML-based distributions-. These novel distributions are formulated by generalizing the exponential function in the canonical BE and FD distributions to the ML function, thereby introducing the tunable parameter $\alpha$ into the quantum statistical weights. This extension preserves the essential quantum characteristics of bosons and fermions while allowing for a more flexible description of particle interactions and thermodynamic behavior, particularly in systems that deviate from idealized equilibrium conditions.
@@
To investigate the physical implications of these generalized distributions, we employ the framework of thermodynamic geometry, which provides a powerful geometric interpretation of thermodynamic state space. By analyzing the thermodynamic curvature and associated metrics, this approach reveals information about microscopic interactions, stability, and phase transitions within the system. Thermodynamic geometry has previously provided important information on systems governed by generalized statistics such as Tsallis, Kaniadakis, and superstatistics [sagi2012observation, saporta2019self, colbrook2017scaling, pessoa2021information, mirza2009nonperturbative, mirza2010thermodynamic, oshima1999riemann], and its application to the MLBE and MLFD distributions provides detailed information on their thermodynamic behavior, including heat capacity and condensation phenomena.
@@ II Broader Context of Generalized Statistics and the Role of Mittag-Leffler Distributions
where $t$ denotes time, $\tau$ is a characteristic relaxation time, and $\beta$ ($0<\beta<1$) is the stretching exponent that quantifies deviations from Debye’s exponential law. At long times, such stretched exponentials asymptotically cross over into power-law decay [bouchaud1992weak]. This crossover is precisely captured by the one-parameter ML function $E_{\alpha}(-t^{\alpha})$, where $\alpha$ ($0<\alpha\leq 1$) is the fractional order parameter: for $\alpha=1$ the function reduces to a pure exponential, while for $0<\alpha<1$ it exhibits algebraic tails.
@@
where $x$ is the displacement of a diffusing particle and the exponent $\alpha$ quantifies the deviation from normal Brownian diffusion (with $\alpha=1$ corresponding to ordinary diffusion, $\alpha<1$ to subdiffusion, and $\alpha>1$ to superdiffusion). Such scaling laws emerge directly from fractional Fokker–Planck equations whose unique solutions are expressed in terms of ML functions [metzler2000random, sandev2011fractional].
@@
where $\mathcal{L}$ denotes the Laplace transform, $s$ is the Laplace variable conjugate to time $t$, and $\lambda>0$ is a characteristic rate parameter. This functional form arises naturally from fractional operators and cannot be represented by ordinary exponential kernels [mainardi2020mittag].
These examples illustrate that the ML function is far more than a mathematical curiosity: it is central to the physics of systems with long memory, anomalous transport, and hierarchical energy landscapes. Indeed, ML functions have long been recognized as the “queen function of fractional calculus,” emerging naturally as the fundamental solution of fractional-order differential equations that describe anomalous diffusion, non-exponential relaxation, and long-memory processes [gorenflo2020mittag, haubold2011mittag]. Its relevance spans numerous areas of physics: modeling viscoelastic relaxation via the fractional Zener model, where stress $\sigma(t)$ and strain $\varepsilon(t)$ are related by fractional derivatives [pritz2003five, nonnenmacher1991fractional]; describing anomalous transport and dispersion in porous or plasma environments via fractional advection–dispersion equations [wheatcraft2008fractional, metzler2000random]; and characterizing non-Debye dielectric relaxation using the Prabhakar (three-parameter ML) function [mainardi2015complete]. Additionally, ML functions provide analytic solutions to fractional kinetic and reaction–diffusion equations relevant in astrophysics, space sciences, and time-series analysis [jose2009generalized, saxena2004generalized].
@@ III Mittag-Leffler Function As a Generalized Distribution
where $\Gamma$ is Gamma function and $\alpha>0$. In the limit $\alpha=1$, the function reduces to the standard exponential $E_{1}(X)=e^{X}$. Moreover, when $\alpha\gg 1$, $E_{\alpha}(X)$ asymptotically approaches 1, reflecting a uniform limiting behavior. The one-parameter form $E_{\alpha}(X)$ seamlessly interpolates between exponential and power-law behaviors. Further generalizations to multi-parameter versions, such as $E_{\alpha,\beta}(X)$ and $E_{\alpha,\beta}^{\gamma}(X)$, enhance this flexibility, enabling more accurate modeling of memory effects and anomalous diffusion in complex systems [mittag1903nouvelle, haubold2011mittag].
Having outlined its mathematical foundation, we now recall the application of the ML function in classical statistical mechanics, where it has been used to generalize the MB distribution  [seifi2025intrinsic]. We introduced the MLMB distribution by generalizing the exponential function in the classical MB distribution to the ML function. This modification enabled a smooth interpolation between classic and quantum (fermion-like and boson-like) behaviors, effectively capturing interaction-induced deviations from classical statistics. In particular, the MLMB distribution exhibits a transition between the classical behavior and the fermion-like (or boson-like) behavior depending on the value of the parameter $\alpha$. When $0<\alpha<1$, the system displays fermionic traits, including negative thermodynamic curvature indicative of repulsive interactions similar to the Pauli exclusion principle. For $1<\alpha\leq 1.5$, it shows bosonic characteristics, with attractive interactions signaling positive curvature and the emergence of Bose-Einstein condensation as the fugacity approaches a critical value. This work demonstrated the versatility of the ML function in extending classical statistical mechanics beyond standard assumptions.
@@
In the MLMB classical distribution [seifi2025intrinsic], the deformation parameter $\alpha$ governs these effective interactions:
@@
This approach parallels other generalized statistical frameworks. For instance, the Tsallis entropic parameter $q$ and the Kaniadakis deformation parameter $\kappa$ [adli2019nonperturbative, mehri2020thermodynamic]similarly tune effective interaction strength and the sign of thermodynamic curvature. The MLMB distribution extends these ideas by enabling a continuous transition between fermion-like (repulsive) and boson-like (attractive) regimes through the single parameter $\alpha$.
@@
As shown in Fig.  1, the MLFD distribution yields an occupation number that remains positive for all values of $\alpha$, approaching unity as $X\to$-$\infty$ and vanishing as $X\to+\infty$. It is noteworthy that for the MLFD distribution with $\alpha>1$, the distribution exhibits anomalous behavior: in the negative domain of $X$, it temporarily exceeds unity before asymptotically approaching 1.
In contrast, Fig. 2 shows that while the MLBE distribution also satisfies $n(X)\to 0$ as $X\to+\infty$, it exhibits divergence near $X=0$ and produces negative values for $X<0$. These negative values of distribution are unphysical, as the occupation number must remain non-negative to ensure physical consistency. The MLFD distribution satisfies this criterion for all $X$ and the entire range of $\alpha$. However, the MLBE distribution imposes a constraint on the fugacity, requiring it to remain within the range $0\leq z\leq 1$ to exclude unphysical results. At the critical value $1<z$, the function $n(X)$ diverges, reflecting the characteristic behavior of the standard BE distribution, where BE condensation occurs [griffin1996bose, zare2012condensation, mirza2011condensation].
@@ V Thermodynamic Curvature of the MLBE and MLFD statistics
Thermodynamic geometry, as developed notably by Ruppeiner and Weinhold, provides a differential geometric structure to the space of thermodynamic parameters [ruppeiner1979thermodynamics, weinhold1975metric]. Within this framework, the space of equilibrium states is treated as a Riemannian manifold, where thermodynamic fluctuations and interactions can be examined through the properties of a metric tensor. The Ruppeiner metric is derived by taking the negative second derivative of entropy with respect to extensive thermodynamic variables, such as internal energy, volume, and particle number. Alternatively, in the energy representation proposed by Weinhold, the metric is obtained from the Hessian of the internal energy with respect to its natural variables.
Applying Legendre transformations to fundamental thermodynamic potentials enables us to construct the equivalent metric structures based on various potentials, such as the Helmholtz and Gibbs free energies. Furthermore, the Fisher information metric—another significant geometric tool—can be introduced through the second derivatives of the logarithm of the partition function with respect to the intensive parameters [ruppeiner1995riemannian, janyszek1990riemannian]:
@@
In systems characterized by two thermodynamic degrees of freedom, the associated geometry is two-dimensional. Janyszek and Mrugała demonstrated that when a metric is derived directly from a thermodynamic potential, the scalar curvature of the parameters space can be expressed in terms of the potential’s second and third derivatives [janyszek1989riemannian]. In two dimensions, the Ricci scalar curvature is computed via:
@@
For an ideal classical gas described by the MB statistics, the thermodynamic curvature vanishes. In contrast, quantum ideal gases exhibit non-zero curvature: it is positive for BE statistics and negative for FD statistics. These signs reflect effective statistical interactions—attractive in the case of bosons and repulsive for fermions [ruppeiner1995riemannian, janyszek1990riemannian, oshima1999riemann]. The divergence of curvature near a critical fugacity (e.g., $z\to 1$ for BE) can signal the presence of phase transitions such as Bose-Einstein condensation. The theoretical framework of geometric thermodynamics has been addressed in earlier studies [mohammadzadeh2017thermodynamic, mohammadzadeh2016perturbative, mirza2011condensation, mehri2020thermodynamic, esmaili2024thermodynamic], which provide comprehensive reviews of its geometric structure and thermodynamic implications.
@@
In the MLBE distribution, for $\alpha<1$, the system displays a fugacity-dependent crossover between fermionic and bosonic characteristics. At a specific point $z=z^{*}_{\alpha}$, indicated by a solid dot in Fig. 3, the thermodynamic curvature changes sign. For fugacity values less than $z^{*}_{\alpha}$, the system displays fermionic characteristics, marked by effective repulsive interactions. In contrast, for $z>z^{*}_{\alpha}$, the system transitions into the bosonic regime, characterized by attractive effective interactions between particles. This crossover, along with the dependence of $z^{*}_{\alpha}$ on the parameter $\alpha$, is illustrated in Fig. 3.
In contrast, when $\alpha>1$, the MLBE distribution exhibits predominantly bosonic behavior for all values of $z$, as indicated by the consistently positive thermodynamic curvature shown in Fig. 4. Moreover, irrespective of the value of $\alpha$, the MLBE distribution features a critical point at $z^{\alpha}_{c}=z^{\alpha}_{c}=1$, where the thermodynamic curvature diverges. This divergence signals the onset of a phase transition analogous to Bose-Einstein condensation, as discussed in Refs. [janyszek1990riemannian, mirza2011condensation]. The critical fugacity $z_{c}$ thus marks the emergence of macroscopic quantum phenomena associated with condensation.
@@
Turning to the MLFD distribution, a similar classification emerges based on $\alpha$. For $\alpha<1$, the system exhibits purely fermionic behavior: the thermodynamic curvature remains negative across all values of $z$, starting from strongly negative values at small $z$ and becoming less negative as $z$ increases, see Fig. (5). In contrast, for $\alpha>1$, the thermodynamic curvature $R$ initially assumes positive values at very small values of $z$. At a specific threshold $z=z^{*}_{\alpha}$, as summarized in the accompanying table in Fig. 6, the thermodynamic curvature undergoes a sign change. Beyond this point, $R$ rapidly transitions to negative values and gradually stabilizes with a mildly decreasing slope as $z$ increases, as shown in Fig. 6.